# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

### v1

@@ 3.1. Deviation Inequalities for Smoothed Plug-in Estimators
The proof of Theorem 3 (see Section 4.3) is similar to that of Theorem 1, and utilizes the variational expression given in (2) to upper bound the desired deviation in terms of empirical process suprema. Note that existing exponential concentration inequalities such as those given in [59] do not apply here since the underlying densities are neither compactly supported nor bounded away from zero.
@@
By joint convexity of Rényi divergences in both its arguments for $\alpha\in[0,1]$ [76, Theorem 11], Theorem 3 also holds with $|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ replaced by $\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$ for $\alpha\in(0,1]$. Similarly, for all $\alpha\in(0,1)\cup(1,\infty)$, $|\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ can be replaced by $\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$ due to convexity of Rényi divergences in the second argument for all orders [76, Theorem 12]. Further, (13) also applies when $\hat{\mu}_{n}*\gamma_{\sigma}$ or $\hat{\nu}_{n}*\gamma_{\sigma}$ (or both) are set to $\mu*\gamma_{\sigma}$ and $\nu*\gamma_{\sigma}$, respectively.
@@
$\tilde{\mathcal{P}}_{M,r,s,\rho}$ $:=\left\{(\mu,\nu)\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}\cap\mathcal{P}_{M,\rho}:p_{\mu},p_{\nu}\in\mathsf{Lip}_{s,1,M}(\rho)\right\}.$
With this, we have the following corollary to Theorem 1.
@@
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\\ \tilde{\mathcal{P}}_{M,r,s,\rho}\end{subarray}}\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)\right|\gtrsim\left(\bar{c}_{d,r}+c_{d,s}M^{2}+(r^{2}+4r)(1+r\sqrt{d})z\right)n^{\frac{-s}{d+2s+4}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq e^{-\left(n^{\frac{d}{d+2s+4}}z^{2}\wedge n^{\frac{d+s+2}{d+2s+4}}z\right)},$ (14)
@@
The proof of Corollary 1 (see Section 4.4) relies on the fact that the difference between smoothed Rényi divergence and its unsmoothed value can be upper bounded given the densities $p_{\mu}$ and $p_{\nu}$ belong to an appropriate Lipschitz class with bounded ratios. The corresponding approximation error is quantified in the next proposition which is a generalization of [67, Lemma 3].
@@
Before we end this subsection, we briefly mention that the deviation inequalities given in Theorem 1 and Theorem 3 could be useful in other contexts in information theory such as physical layer security or covert communications, where the secrecy metric of interest is often expressed in terms of Rényi divergences. Exponential deviation inequalities then provide high confidence bounds in random coding- or expurgation-based arguments to show existence of codebooks satisfying the security criteria approximately. For instance, consider the case of information-theoretically secure coding over a Gaussian wiretap channel with a per-codeword power constraint. In this context, $\{X_{i}\}_{i=1}^{n}$ would denote a random codebook of size $n$ with each codeword of length $d$ sampled independently according to $\mu$ and $\hat{\mu}_{n}*\gamma_{\sigma}$ would correspond to the output distribution of a discrete-time additive Gaussian noise channel of variance $\sigma^{2}$ with input uniformly sampled from the codebook. Then, it is of interest to show that $\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\leq\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)+o(1)$ with overwhelming probability, where $\nu$ is an auxiliary distribution of interest. To the best of our knowledge, existing deviation inequalities relevant to this setting are applicable only to bounded metrics or finite (output-) alphabet channels (see e.g., [44, Theorem 31] and [25, Lemma 4]). Hence, we expect Theorem 1 and 3 to have potential new applications in wiretap coding or covert communications over Gaussian channels.
@@ 3.2. Deviation Inequalities for Neural Estimators of Rényi Divergences
Let $\beta>d/2$, $M\geq 1$, $a:=\frac{d}{2\beta}$, $b:=\log M$, $\mathcal{G}_{b,\beta,\rho}$ be a neural class satisfying Assumption 2, and $c_{d,\beta}$ be such that $1\leq c_{d,\beta}\lesssim\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}$, where $\hat{c}_{d,\beta}$ is as given in (17). Then, the following hold:
@@
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq 2e^{-(nz^{2}\wedge nz)}.$ (18)
@@
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq 2e^{-(nz^{2}\wedge nz)},$ (19)
@@
3.3. Application: Auditing Rényi DP
@@
As $\epsilon$-DP is equivalent to $\sup_{u\sim v}\mathsf{D}_{\max}(\mu_{u}\|\mu_{v})\leq\epsilon$, where $\mathsf{D}_{\max}$ is the $\infty$-order Rényi divergence and $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$ is non-decreasing in $\alpha$ [76, Theorem 3], it is clear that Rényi DP of any finite order $\alpha$ is a relaxation of DP. However, the two notions become equivalent at $\alpha=\infty$, and hence $\epsilon$- Rényi DP is a good approximation of $\epsilon$-DP for sufficiently large $\alpha$.
@@
$\begin{split}&H_{0}:\mathsf{D}_{\alpha}\left(\mu_{0}\middle\|\nu_{0}\right)\leq\epsilon,\\ &H_{1}:\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)>\epsilon.\end{split}$ (20)
@@
$e_{1,n}(T_{n},t_{n})$ $\leq 2e^{-n^{\tau}},$
$e_{2,n}(T_{n},t_{n})$ $\leq 2e^{-\vartheta_{n,\alpha,\delta_{n}}(\tau)},$
@@ 4. Proofs
$\mathcal{F}_{\sigma}(\mathcal{F}):=\big{\{}f*\varphi_{\sigma}:f\in\mathcal{F}\big{\}}\quad\mbox{and }\quad\mathcal{E}_{\sigma,\alpha}(\mathcal{F}):=\big{\{}e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}\big{\}},$ (21)
be the Gaussian-smoothed function classes obtained by convolving the functions in $\mathcal{F}$ and $e^{\alpha\mathcal{F}}:=\{e^{\alpha f}:f\in\mathcal{F}\}$ by $\varphi_{\sigma}$, respectively. We will denote $\mathcal{E}_{\sigma,1}(\mathcal{F})$ by $\mathcal{E}_{\sigma}(\mathcal{F})$ henceforth.
@@ 4.1. Proof of Theorem 1
We require the following lemma which provides a general deviation inequality for Gaussian-smoothed KL divergence.
@@
Let $\sigma>0$, $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, $X^{n}\sim\mu^{\otimes n}$ and $Y^{n}\sim\nu^{\otimes n}$. Suppose $\mathcal{F}_{n}$ with envelope $F_{n}$ is such that $\mathcal{F}_{n,\sigma}:=\mathcal{F}_{\sigma}(\mathcal{F}_{n})$ and $\mathcal{E}_{n,\sigma}:=\mathcal{E}_{\sigma}(\mathcal{F}_{n})$ with envelopes [^7] By Young’s convolution inequality, a sufficient condition for existence of a measurable envelope $F_{n}*\varphi_{\sigma}$ for $\mathcal{F}_{n,\sigma}$ is $F_{n}\in L^{p}(\mathbb{R}^{d})$ for some $p\geq 1$. $F_{n}*\varphi_{\sigma}$ and $e^{F_{n}}*\varphi_{\sigma}$, respectively, satisfy Assumption 1. Further, assume that $f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star},f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\in\mathcal{F}_{n}$ a.s., and $F_{n}$ satisfies
$\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}\vee\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}<\infty,$ (23)
for some $q\in(0,1]$ and every $n\in\mathbb{N}$, where $X\sim\mu$ and $Y\sim\nu$. Then
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\gtrsim_{q}n^{-\frac{1}{2}}t_{n,\sigma,\mu,\nu}+z\right)$
$\qquad\qquad\qquad\qquad\leq e^{-R_{n,q,\sigma,\mu}(z)}+e^{-\tilde{R}_{n,q,\sigma,\nu}(z)}+3e^{-O_{n,q,\sigma,\mu}(z)}+3e^{-\tilde{O}_{n,q,\sigma,\nu}(z)},$ (24)
@@
$t_{n,\sigma,\mu,\nu}:=J(1,F_{n}*\varphi_{\sigma},\mathcal{F}_{n,\sigma},\mu)\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,e^{F_{n}}*\varphi_{\sigma},\mathcal{E}_{n,\sigma},\nu)\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\nu},$ (25a)
[…]
$O_{n,q,\sigma,\mu}(z):=\frac{n^{q}z^{q}}{\left\|M_{n,\mathcal{F}_{n,\sigma}}(X^{n})\right\|_{\psi_{q}}^{q}}\quad\mbox{ and }\quad\tilde{O}_{n,q,\sigma,\nu}(z):=\frac{n^{q}z^{q}}{\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}(Y^{n})\right\|_{\psi_{q}}^{q}},$ (25c)
$M_{n,\mathcal{F}_{n,\sigma}}(X^{n}):=\max_{1\leq i\leq n}F_{n}*\varphi_{\sigma}(X_{i})\quad\mbox{ and }\quad\tilde{M}_{n,\mathcal{E}_{n,\sigma}}(Y^{n}):=\max_{1\leq i\leq n}e^{F_{n}}*\varphi_{\sigma}(Y_{i}),$ (25d)
[…]
$\check{M}_{n,\mathcal{E}_{n,\sigma}}\vee\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}(Y^{n})\right\|_{\psi_{q}}\lesssim_{q}\left(\log(1+n)\right)^{\frac{1}{q}}\left\|e^{F}_{n}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}<\infty.$ (25g)
When $\mu,\nu$ are compactly supported and (23) is replaced by the condition
@@
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\gtrsim n^{-\frac{1}{2}}t_{n,\sigma,\mu,\nu}+z\right)$
@@
The proof of Lemma 2 is given in Section 5.1. The key idea in its proof is to use the variational expression (1b) to upper bound the deviation $|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ in terms of sum of suprema of empirical processes indexed by $\mathcal{F}_{n,\sigma}$ and $\mathcal{E}_{n,\sigma}$. Then, (24) and (26) follow by invoking a Fuk-Nagaev type inequality given in [2, Theorem 4] (see Theorem 7 below) and Talagrand’s concentration inequality given in [12, Theorem 2.3] (see Theorem 6 below), respectively.
@@
From the proof of Lemma 2, it is easy to see that (24) and (26) holds with $\nu*\gamma_{\sigma}$ and $\hat{\nu}_{n}*\gamma_{\sigma}$ replaced with $\nu$ such that $\mu*\gamma_{\sigma},\hat{\mu}_{n}*\gamma_{\sigma}\ll\nu$ if the integrability conditions on only $F_{n}$ hold, i.e., $\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}<\infty$ and $\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}<\infty$, respectively. For instance, (24) simplifies to
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu\right)\right|\gtrsim_{q}n^{-\frac{1}{2}}J(1,F_{n}*\varphi_{\sigma},\mathcal{F}_{n,\sigma},\mu)\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}+z\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-R_{n,q,\sigma,\mu}(z)}+3e^{-O_{n,q,\sigma,\mu}(z)}.$ (27)

### v3

@@ 3.1. Deviation Inequalities for Smoothed Plug-in Estimators
The high-level idea used in the proof of Theorem 3 (see Section 4.3) is similar to that of Theorem 1, and utilizes the variational expression given in (2) to upper bound the desired deviation in terms of empirical process suprema. However, the presence of logarithms in the expression necessitates a linearization step prior to achieving this goal. Fortunately, this is possible when $\mu,\nu$ are compactly supported by noting that the Lipschitz constant of logarithm is bounded by $c$ on the interval $[1/c,c]$ for $c\in[1,\infty)$. Notice that in contrast to Theorem 1, we do not have the counterpart of (9) for sub-Gaussian $\mu,\nu$. This is because even in the one-sample setting, the truncation argument, required for similar reasons as in Part $(ii)$ of Theorem 1, results in the norm of the envelope of the function class indexing the empirical process suprema, scaling exponentially with respect to the truncation level.
@@
By joint convexity of Rényi divergences in both its arguments for $\alpha\in[0,1]$ [84, Theorem 11], Theorem 3 also holds with $|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ replaced by $\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$ for $\alpha\in(0,1]$. Similarly, for all $\alpha\in(0,1)\cup(1,\infty)$, $|\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ can be replaced by $\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$ due to convexity of Rényi divergences in the second argument for all orders [84, Theorem 12]. Further, (15) also applies when $\hat{\mu}_{n}*\gamma_{\sigma}$ or $\hat{\nu}_{n}*\gamma_{\sigma}$ (or both) are set to $\mu*\gamma_{\sigma}$ and $\nu*\gamma_{\sigma}$, respectively.
@@
$\tilde{\mathcal{P}}_{M,r,s,\rho}:=\left\{(\mu,\nu)\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}\cap\mathcal{P}_{M,\rho}:p_{\mu},p_{\nu}\in\mathsf{Lip}_{s,1,M}(\rho)\right\},$
where $\mu\ll\gg\nu$ means that $\mu\ll\nu$ and $\nu\ll\mu$. With this, we have the following corollary to Theorem 1.
@@
[…]
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-\left(n^{\frac{d}{d+2s+4}}z^{2}\wedge n^{\frac{d+s+2}{d+2s+4}}z\right)},$ (16)
@@
The proof of Corollary 1 (see Section 4.4) relies on the fact that the difference between smoothed Rényi divergence and its unsmoothed value can be upper bounded given the densities $p_{\mu}$ and $p_{\nu}$ belong to an appropriate Lipschitz class with bounded ratios. The corresponding approximation error is quantified in the next proposition which is a generalization of [76, Lemma 3].
@@
Before we end this subsection, we briefly mention that the deviation inequalities of the form given in Theorem 1 and Theorem 3 could be useful in other contexts in information theory such as physical-layer security or covert communications, where the secrecy metric of interest is often expressed in terms of Rényi divergences. Exponential deviation inequalities then provide high confidence bounds in random coding- or expurgation-based arguments to show existence of codebooks satisfying the security criteria approximately. For instance, consider the case of information-theoretically secure coding over a Gaussian wiretap channel with a per-codeword power constraint. In this context, $\{X_{i}\}_{i=1}^{n}$ would denote a random codebook of size $n$ with each codeword of length $d$ sampled independently according to $\mu$, and $\hat{\mu}_{n}*\gamma_{\sigma}$ would correspond to the output distribution of a discrete-time additive white Gaussian noise channel of variance $\sigma^{2}$ with input uniformly sampled from the codebook. Then, it is of interest to show that $\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\leq\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)+o(1)$ with overwhelming probability, where $\nu$ is an auxiliary distribution of interest. To the best of our knowledge, existing deviation inequalities relevant to this setting are applicable only to bounded metrics or finite (output-) alphabet channels (see e.g., [54, Theorem 31] and [31, Lemma 4]). Hence, we expect Theorem 1 and 3 to have potential new applications in wiretap coding or covert communications over Gaussian channels when sufficient shared randomness is available. For wider applicability, it appears that an improvement of the dependence of dimension in these bounds would be required.
@@ 3.2. Deviation Inequalities for Neural Estimators of Rényi Divergences
Let $\beta>d/2$, $M\geq 1$, $a:=d/(2\beta)$, $b:=\log M$, $\mathcal{G}_{b,\beta,\rho}$ be a neural class satisfying Assumption 2, and $c_{d,\beta}$ be such that $1\leq c_{d,\beta}\lesssim\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}$, where $\hat{c}_{d,\beta}$ is as given in (19). Then, the following hold:
@@
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 2e^{-(nz^{2}\wedge nz)}.$ (20)
@@
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 2e^{-(nz^{2}\wedge nz)},$ (21)
@@
3.3. Application: Auditing Rényi Differential Privacy
@@
As $\epsilon$-DP is equivalent to $\sup_{u\sim v}\mathsf{D}_{\max}(\mu_{u}\|\mu_{v})\leq\epsilon$, where $\mathsf{D}_{\max}$ is the $\infty$-order Rényi divergence and $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$ is non-decreasing in $\alpha$ [84, Theorem 3], it is clear that Rényi DP of any finite order $\alpha$ is a relaxation of DP. However, the two notions become equivalent at $\alpha=\infty$, and hence $\epsilon$- Rényi DP is a good approximation of $\epsilon$-DP for sufficiently large $\alpha$.
@@
$H_{0}:\mathsf{D}_{\alpha}\left(\mu_{0}\middle\|\nu_{0}\right)\leq\epsilon\quad\text{against}\quad H_{1}:\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)>\epsilon.$ (22)
@@
$e_{1,n}(T_{n},t_{n})\leq 2e^{-n^{\tau}}\quad\text{and}\quad e_{2,n}(T_{n},t_{n})\leq 2e^{-\vartheta_{n,\alpha,\delta_{n}}(\tau)},$
@@ 4. Proofs
$\mathcal{F}_{\varphi_{\sigma}}(\mathcal{F}):=\big{\{}f*\varphi_{\sigma}:f\in\mathcal{F}\big{\}}\quad\mbox{and }\quad\mathcal{E}_{\varphi_{\sigma},\alpha}(\mathcal{F}):=\big{\{}e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}\big{\}},$ (23)
be the Gaussian-smoothed function classes obtained by convolving the functions in $\mathcal{F}$ and $e^{\alpha\mathcal{F}}:=\{e^{\alpha f}:f\in\mathcal{F}\}$ by $\varphi_{\sigma}$, respectively. We will denote $\mathcal{E}_{\varphi_{\sigma},1}(\mathcal{F})$ by $\mathcal{E}_{\varphi_{\sigma}}(\mathcal{F})$ henceforth.
@@ 4.1. Proof of Theorem 1
We will state a more general lemma than required for our purposes, which provides a deviation inequality for smoothed plug-in estimators of KL divergence satisfying certain regularity conditions. Let $\Phi\in\mathcal{P}(\mathbb{R}^{d})$, and $\phi$ denote its Lebesgue density which will act as a smoothing kernel. Note that $\left\|\phi\right\|_{1,\mathbb{R}^{d}}=1$. Set $\mathcal{F}_{\phi}(\mathcal{F}):=\big{\{}f*\phi:f\in\mathcal{F}\big{\}}$ and $\mathcal{E}_{\phi}(\mathcal{F}):=\big{\{}e^{f}*\phi:f\in\mathcal{F}\big{\}}$.
@@
Let $\pi\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ with marginals $\mu$ and $\nu$ be such that $\mu\ll\nu$. Suppose $\mathcal{F}_{n}$ with envelope $F_{n}$ is such that $\mathcal{F}_{n,\phi}:=\mathcal{F}_{\phi}(\mathcal{F}_{n})$ and $\mathcal{E}_{n,\phi}:=\mathcal{E}_{\phi}(\mathcal{F}_{n})$ with envelopes [^7] By Young’s convolution inequality, a sufficient condition for existence of a measurable envelope $F_{n}*\phi$ for $\mathcal{F}_{n,\phi}$ is $F_{n}\in L^{p}(\mathbb{R}^{d})$ for some $p\geq 1$. $F_{n}*\phi$ and $e^{F_{n}}*\phi$, respectively, satisfy Assumption 1, $f_{\mu*\Phi,\nu*\Phi}^{\star}\in\mathcal{F}_{n}$, $F_{n}*\phi\in L^{p}(\mu)$ and $e^{F_{n}}*\phi\in L^{p}(\nu)$ for some $p\geq 2$. Further, assume that for $(X^{n},Y^{n})\sim\pi^{\otimes n}$, $\hat{\mu}_{n}*\Phi\ll\hat{\nu}_{n}*\Phi$ and $f_{\hat{\mu}_{n}*\Phi,\hat{\nu}_{n}*\Phi}^{\star}\in\mathcal{F}_{n}$ a.s. Then
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\hat{\nu}_{n}*\Phi\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)\right|\gtrsim_{p}n^{-\frac{1}{2}}t_{n,\phi,\mu,\nu}+z\right)$
$\qquad\qquad\qquad\leq e^{-R_{n,p,\phi,\mu}(z)}+e^{-\tilde{R}_{n,p,\phi,\nu}(z)}+\Big{(}\left\|F_{n}*\phi\right\|_{p,\mu}^{p}+\left\|e^{F_{n}}*\phi\right\|_{p,\nu}^{p}\Big{)}n^{1-p}z^{-p},$ (25)
@@
$t_{n,\phi,\mu,\nu}:=J(1,F_{n}*\phi,\mathcal{F}_{n,\phi},\mu)\left\|F_{n}*\phi\right\|_{2,\mu}+J(1,e^{F_{n}}*\phi,\mathcal{E}_{n,\phi},\nu)\left\|e^{F_{n}}*\phi\right\|_{2,\nu},$ (26a)
$R_{n,p,\phi,\mu}(z):=\frac{nz^{2}}{\left\|F_{n}*\phi\right\|_{2,\mu}^{2}}\wedge\frac{n^{\frac{p-1}{p}}z}{\left\|F_{n}*\phi\right\|_{p,\mu}}\quad\mbox{and}\quad\tilde{R}_{n,p,\phi,\nu}(z):=\frac{nz^{2}}{\left\|e^{F_{n}}*\phi\right\|_{2,\nu}^{2}}\wedge\frac{n^{\frac{p-1}{p}}z}{\left\|e^{F_{n}}*\phi\right\|_{p,\nu}}.$ (26b)
In particular, when $\mu,\nu$ are compactly supported, $\phi=\varphi_{\sigma}$ and
@@
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\gtrsim n^{-\frac{1}{2}}t_{n,\varphi_{\sigma},\mu,\nu}+z\right)$
@@
The proof of Lemma 2 is given in Section 5.1. The key idea in its proof is to use the variational expression (1b) to upper bound the deviation $|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\hat{\nu}_{n}*\Phi\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)|$ in terms of sum of suprema of empirical processes indexed by $\mathcal{F}_{n,\phi}$ and $\mathcal{E}_{n,\phi}$. Then, (25) and (27) follow by invoking a Fuk-Nagaev type inequality given in [3, Theorem 2] (see Theorem 7 below) and Talagrand’s concentration inequality given in [15, Theorem 2.3] (see Theorem 6 below), respectively. While (27) will suffice for our purposes, Lemma 2 is more general both in terms of the flexibility of the smoothing kernel $\phi$ as well as relaxed regularity conditions on the probability distributions. For instance, (25) may be used to derive deviation inequalities in the two-sample Gaussian-smoothed setting when $\left\|X-Y\right\|_{2}$ is bounded a.s. for $(X,Y)\sim\pi$ and the tails of the marginals $\mu,\nu$ decay faster than sub-Gaussian. By following similar steps as in the proof below, it is easy to see that this case fits within the framework of Lemma 2 since the aforementioned boundedness and tail decay condition ensures that $f_{\hat{\mu}_{n}*\varphi,\hat{\nu}_{n}*\varphi}^{\star},f_{\mu*\varphi,\nu*\varphi}^{\star}\in\mathcal{F}_{n}$ for an $\mathcal{F}_{n}$ with envelope $F_{n}(x)=O_{\pi}(1+\left\|x\right\|^{2})$ such that the exponential moments of $F_{n}*\varphi$ exists.
@@
From the proof of Lemma 2, it is easy to see that (25) and (27) holds with $\nu*\Phi$ and $\hat{\nu}_{n}*\Phi$ replaced with $\nu$ such that $\mu*\Phi,\hat{\mu}_{n}*\Phi\ll\nu$ if the integrability conditions on only $F_{n}$ hold, i.e., $\left\|F_{n}*\phi\right\|_{p,\mu}<\infty$ and $\left\|F_{n}*\phi\right\|_{2,\mu}\vee\left\|F_{n}*\phi\right\|_{\infty,\mu}<\infty$, respectively. For instance, (25) simplifies to
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\nu\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu\right)\right|\gtrsim_{q}n^{-\frac{1}{2}}J(1,F_{n}*\phi,\mathcal{F}_{n,\phi},\mu)\left\|F_{n}*\phi\right\|_{2,\mu}+z\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-R_{n,p,\phi,\mu}(z)}+\left\|F_{n}*\phi\right\|_{p,\mu}^{p}n^{1-p}z^{-p}.$